OpenStudy (anonymous):

\[\frac{ 7x-21 }{ 18-6x }\]

hero (hero):

Hint: \[\frac{ 7x-21 }{ 18-6x }=-\frac{7x - 21}{6x - 18}\]

OpenStudy (anonymous):

x-3

hero (hero):

\[\frac{ 7x-21 }{ 18-6x }=-\frac{7x - 21}{6x - 18} = -\frac{7(x-3)}{6(x-3)}\]

hero (hero):

Cancel common factors to get the final result.

OpenStudy (anonymous):

ooooo i get you now

OpenStudy (anonymous):

when are you going to reach 100

hero (hero):

Never

OpenStudy (anonymous):

lol y not

hero (hero):

Because no one will overtake @Satellite73

OpenStudy (anonymous):

what you mean over take

hero (hero):

In order to get 100, you have to beat @Satellite's Smart Score. Even if Satellite died, it would still take five years to catch up with him.

OpenStudy (anonymous):

\[\frac{ 5z }{ \sqrt{z^5} }\]

hero (hero):

Rationalize the denominator

OpenStudy (anonymous):

o wow and 5/x^4

OpenStudy (anonymous):

sooory nvm

OpenStudy (anonymous):